INVESTIGATION OF J-DEPENDANCE OF LINE SHIFT AND LINE BROADENING IN THE ν_{1}+3ν_{3} BAND OF ACETYLENE

H. VALIPOUR and \underline{D. ZIMMERMANN}, Fachbereich Physik, Technische Universität Berlin, Germany.

Line shift and line broadening coefficients δν and γ have been measured for 21 rotational lines of the P and R branch of the ν_{1}+3ν_{3} band of acetylene at 789 nm using a tunable cw semiconductor laserdiode system. The absorption cell was of the multi-pass Herriot-type with an effective absorption length of 35 m. Up to now, self-broadening and broadening by N_{2}, Ar and air perturbers have been investigated. Taking air-broadening as an example we obtained γ = 0.176(12) cm^{−1}/bar and δν = 0.012(2)cm^{−1}/bar for the P(4) line. With increasing value of J the values of γ decrease, e.g. to 0.147(2) cm^{−1}/bar for R(13) whereas δν increases to 0.016(2) cm^{−1}/bar. A similar behaviour has been observed for the case of other foreign gas perturbers and for self-broadening. The present investigation might be of interest for the reliable detction of small amounts of acetylene in air of atmospheric pressure. In our system we could detect a minimal concentration of 1 μbar acetylene in a N_{2} atmosphere of 15 mbar pressure.